(8c6p85) A car weighing 1.07×104 N and traveling at 14.00 m/s attempts to round an unbanked curve with a radius of 51 m, what force of friction is required to keep the car on its circular path? Submit Answer Tries 0/7 If the coefficient of static friction between the tires and road is 0.500, is the attempt at taking the curve successful? Submit Answer Tries 0/7
Answer #1

a) The mass of the car is

m = W/g 1.07 × 104/9.8 = 1091.84 kg

The force of friction required is the centripetal force

F = mv2/r = 109 1.84 × 142/51 4196,08

b) the coefficient of friction required is

μ-F/W 4196.08/1.07 × 104 0.39

Since the coefficient of static friction is more than the required values of friction, the attempt is successful. Answer is yes